# Second Order Stein-Variational Dynamic Optimization

### v1

@@ 1 Introduction
Trajectory optimization is essential for robotic and autonomous systems to operate safely in dense and cluttered environments. From an optimization point of view, the planning and trajectory optimization problems in robotics have some unique characteristics. These include the non-convexity of the objective function in consideration, the nonlinear state and actuation constraints, and the highly nonlinear dynamics that characterize most robotic systems. The characteristic of non-convexity is due to the existence of multiple local minima. Such situations arise in reaching and navigation tasks performed in cluttered and dense environments.
There is a long history of trajectory optimization methods and their application to robotics and autonomous systems. Starting from the Differential Dynamic Programming (DDP) [1] and the Iterative Pontryagin Maximum Principle (iPMP) [2], these optimal control methods have been extensively used in robotics. Both DDP and iPMP belong to the same family of algorithms that are gradient-based and rely on forward-backward sweeps. Their mathematical structure is derived using either the Bellman Principle of Optimality or the Pontryagin Maximum Principle and consists of the backpropagation of the value function or costate and forward propagation of state trajectories. This step of backpropagation is performed locally around the trajectories that are updated at each iteration. While the locality of the backward pass makes DDP and iPMP scalable and fast, it often results in sub-optimal solutions.
@@
In this paper, we investigate two mechanisms that combine sampling with gradient-based optimization and parallelism, and develop algorithms for planning and model predictive control with highly non-convex objectives. The first mechanism is a schedule of optimization steps, that consists of number of Netwon steps followed by an iteration that relies on sampling. The motivation for this hybrid mechanism is that while Newton steps can locally optimize the objective function, sampling can disrupt this process to ensure that optimization is not getting trap into a sub-optimal local minima. In the context of trajectory optimization for robotics, this mechanism was first introduced in the work on Maximum Entropy DDP (MEDDP) and its uni-modal and multi-modal extensions. The difference of this work with respect to [17] is that here we show the benefits of this mechanisms in the context of MPC. It is also interesting to note that this hybrid optimization mechanism, which we name Messy Optimization, was also proposed in the area of business administration and management science [18] as mental model for enhances creativity and improving innovation.
The second mechanism goes one step further than MEDDP and results in new algorithm, the Stein Variational Differential Dynamic Programming (SVDDP). Grounded on the Stein Variational Newton’s method [19], SVDDP is a kernel-based extension of MEDDP. In SVDDP the optimization process alternates between two phases. Starting from different initial trajectories/particles, the first phase consists of multiple DDP updates performed in parallel fashion resulting in optimized state and control trajectories. In the second phase, these trajectories are updated only once with DDP using the Stein-Variational Newton to update the controls and generate the new state trajectories. This phase enforces exploration by pushing the trajectories to be stay away from each other. In addition the fact that this update is performed once every few iterations of the first phase ensures that the overall optimization will not be disrupted. We benchmark the proposed algorithm against variations on MEDDP and MPPI for both trajectory optimization and MPC and show that it matches and, in most cases, outperforms these algorithms. The rest of the paper is organized as follows. Section 2 introduces the background of our work which includes MEDDP and SV gradient descent and Newton’s method. In Section 3 we derive SVDDP. Section 4 includes experimental results, and section 5 concludes with future extensions.
@@ 2 Background
2.1 Maximum Entropy Dynamic Second-Order Optimization
@@ 2.2 Maximum Entropy DDP
where $Q(x,u)$ is the state-action value function in DDP and $Z(x)$ is the corresponding partition function. We refer readers to [17] and Appendix A for derivation. This indicates that the optimal policy has a form of unimodal Gaussian $\pi^{\ast}(\delta u|\delta x)\sim\mathcal{N}(\delta u^{\ast},\alpha Q_{uu}^{-1})$, where $\delta u^{\ast}$ is a solution of vanilla deterministic DDP:
$\delta{u}^{\ast}=\kappa+K\delta x,\quad\text{with}\ \kappa=-{Q}^{-1}_{{uu}}{Q_{{u}}},\ K=-{Q}^{-1}_{{uu}}{Q_{{ux}}}.$
@@ 2.3 Stein Variational Gradient Descend and Newton’s method
This section briefly covers Stein Variational Gradient Descent (SVGD) and SV Newton’s Method (SVNM). We first introduce the update law of SVGD, then describe the connection to functional gradient descent. Finally, we review SVNM as an extension to the second-order optimization of SVGD. For further information, we refer the reader to [16] for SVGD and [22] for SVNM. In this section, we use $x\in\mathbb{R}^{d}$ to denote a variable of a function but not a state of a system. In addition, we use $x_{n}\in\mathbb{R}^{d}$ to denote a sampled point.
Stein Variational Gradient Descend: Let $p$ on $\mathbb{R}^{d}$ be a target distribution that we would like to approximate using a collection of samples. We drop the argument of the function $p$, and specify it only when we evaluate it at point $x_{n}$ as $p(x_{n})$. With samples $\{x_{n}\}$ from a tractable reference distribution $q$ on $\mathbb{R}^{d}$, SVGD iteratively computes a transport map $T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ so that the transformed samples from $p$, i.e., $\{T(x_{n})\}$ can be an empirical approximation of $p$. This map is obtained by solving the following optimization problem.
@@
where $l$ stands for $l$ th iteration. Here, $D_{\rm{KL}}$ is KL divergence that measures the difference between the two distributions. To simplify this problem, SVGD considers the vector-valued Reproducing Hilbert Kernel Space (RKHS) $\mathcal{H}^{d}=\mathcal{H}\times\cdots\times\mathcal{H}$, where $\mathcal{H}$ is a scalar-valued RKHS with kernel $k(x,x^{\prime})$. With RKHS, one can represent a function as a weighted composition of kernels, centered at samples. This representation also allows us to perform optimization in functional space by working with the weights in the same way as the normal optimization technique, such as GD. Now, the transportation map is written as a perturbation of the identity map denoted by $I$ along $\mathcal{H}^{d}$ as
@@
Observe the similarity with GD update of a normal optimization, i.e., $x^{l+1}=x^{l}+\epsilon\nabla f_{0}(x^{l})$, with a step size $\epsilon$ and objective $f_{0}$. With this $T^{l}$, we define the objective of the problem (6) as
$\hat{J}_{q^{l}}[Q]=D_{\rm{KL}}\big{(}(I+Q)\ast q^{l}|p\big{)},$ (7)
which shold be decreased by the proper choice of $Q$ from $\hat{J}_{q^{l}}[\bm{0}]$, where $\bm{0}(x)=0$ is a zero map. Since this $Q$ lies in the steepest direction of the functional $\hat{J}_{q}^{l}$ at $\bm{0}$, the map is given as follows.
@@
The descent direction satisfies the following condition with $S\in\mathcal{H}^{d}$,
@@ 3.1 ME DDP with Functional Gradient and Hessian
We consider $N$ trajectories optimized by DDP that we would like to spread out for exploration. Using the deviation $u^{\ast}=\bar{u}+\delta u^{\ast}$, the gradient operator $\nabla_{u\ast}$ for $Q$ is equivalent to $\nabla_{\delta u}$. Therefore, with the quadratically approximated $Q$, and the substitution of $\pi^{\ast}(u^{\ast})$ for $p(x)$ in the functional gradient (9), we get:
@@
where the derivatives of $Q$ are evaluated at $\bar{u}$. The first term on the right-hand side is zero because of the optimality condition used in the derivation of DDP. Therefore, we are left with the repulsive force terms, which seem attractive for spreading trajectories. However, since we lose the effect of $\alpha$, this formulation cannot capture the relative significance of the entropy in the objective. Thus, we need SVNM. By substituting $\pi^{\ast}(u)$ for $p(x)$ in the approximated Hessian (12) with the terms in (11), and performing empirical approximation, the Hessian for SV Newton’s method is obtained as:
$H^{s,s}=\frac{1}{N}\sum_{n=1}^{N}\Big{[}\frac{Q^{(n)}_{uu}}{\alpha}k(u^{(n)},u^{(s)})^{2}+\nabla_{u^{(n)}}k(u^{(n)},u^{(s)})\nabla k_{u^{(n)}}(u^{(n)},u^{(s)})^{\mathsf{T}}\Big{]},$ (14)
@@
$u^{(s)}\leftarrow u^{(s)}+\epsilon w(u^{(s)}),\quad\text{with}\quad w(u^{(s)})=\sum_{n=1}^{N}\beta^{(n)}k(u^{(n)},u^{(s)}).$ (15)
Now, $\alpha$ is captured in $H^{s,s}$. Hence, we believe that this should be the correct formulation to incorporate the SV method with DDP. The second term in (14) is known as an expected outer product of the gradient, which captures information on the variation in all directions [23]. Thus, it can be seen as an approximated Hessian of the kernel. We also provide another example of interpretation in the Appendix B.2. To comprehend how the Hessian affects the repulsive force term (second term of the right hand-side of (13)), let us consider a simple example with $H=Q_{uu}\in\mathbb{R}^{2\times 2}$. We can draw an ellipsoid from its eigenvectors and eigenvalues whose major and minor axes point to the directions where $Q$ increases most rapidly and slowly, respectively. When the matrix is inverted, the scaling factors of the axes, which are the eigenvalues, are also inverted. However, the eigenvectors are preserved. Consequently, the major and minor axes are flipped. Hence, the major and minor axes of $Q_{uu}^{-1}$ indicate the direction in which $Q$ increases most slowly and rapidly, respectively. When (12) is solved, the repulsive force term is multiplied by $Q_{uu}^{-1}$. This multiplication projects repulsive force more in the direction where $Q$ increases slowly. In other words, the repulsive force is transformed so that it does not disturb the optimization process too much, leading to efficient exploration. We provide the schematic of SVDDP in Fig.1. We note that the update (15) is not applied to the best trajectory, ensuring that at least one mode shows a monotonic cost reduction to preserve the convergence property of DDP as in [17, 24]. We note that we use cost as a metric, although other metrics e.g., diversity, can be used depending on the purpose.
@@ 4 Experimental Results
This section provides experimental results, including Trajectory Optimization (TO) and MPC. For TO, we compare normal DDP, Unimodal, Multimodal Gaussian DDPs (UG-, MG-MEDDP), and SVDDP. We note that DDPs are modified so that they can handle constraints as in section 3.3. For MPC, our comparisons involve DDP-MPC, UG-MEDDP, MG-MEDDP, SVDDP in MPC mode. In addition, MPPIs with three different policy parameterization and update laws are included in the comparison. They are Unimodal Gaussian MPPI (UG-MPPI), Multimodal Gaussian MPPI (MG-MPPI), and Unimodal Gaussian with Stein Variational update law (SV-MPPI) [7]. These three MPPIs correspond to the UG-MEDDP, MG-MEDDP, and SVDDP, respectively. In terms of systems, we use three systems, including a 2D Car, quadrotor, and 7DoF arm in TO and 2D car and quadrotor for MPC. In addition, we have comparisons of DDPs in swimmer and ant dynamics in Brax [31] whose results are presented in section F.
Trajectory Optimization: This experiment is to examine the exploration capability and convergence of SV and ME DDPs over normal DDP. In addition, we observe the effect of temperature $\alpha$, number of modes, and stepsize $\epsilon$ on the performance. Figure D.1 provides an overview of the experiments. We performed the same experiment 15 times to capture the properties of the algorithms, and 15 outcomes are shown in the figure. The task is to reach the target while avoiding obstacles under control limits. The configuration and shapes of the obstacles give multiple local minima, which normal DDP without exploration gets captured by poor ones. ME and SVDDPs can explore and find better solutions with the exploration mechanisms. To examine the effect of temperature $\alpha$ and the number of modes (trajectories), we performed the same experiments, with different values of them. The results of the experiments are provided in Table D.1. For SVDDP, we also changed $\epsilon$s. We plot the evolution of the mean cost over iteration in Fig. D.2 to compare the convergence speed of DDPs.
@@
Model Predictive Control: We perform experiments in MPC with the 2D car and quadrotor system used in TO for similar reaching tasks. With 2D car and quadrotor, we compare seven algorithms of MPC DDPs and MPPIs to examine the performance of our algorithms over one of the most successful sampling-based algorithms MPPI. We have DDP, UG-MEDDP, MG-MEDDP, and SVDDP in MPC mode for DDP-based algorithms. For MPPI, we have UG-, MG-, and SV-MPPI. Details of the experiments are provided in Appendix E. We use [32] to tune MPPIs parameters such as covariance and temperature, etc.
@@
In the experiments, a trajectory is considered successful when it reaches a ball centered around the targets with radius 0.5, and does not hit the obstacles. The experiments are performed ten times per algorithm and environment, except for the deterministic normal DDP. This means for each algorithm we will have 100 runs from which we collect statistics on the performance.
@@
2D Car: One of the obstacle fields and results are shown in Fig. 2(a) and Fig. 2(b). The initial point and the target are [0, 0] and [5, 5]. Circular obstacles have a radius of 0.5. Statistics on the performance of the algorithms with respect to the success rate are provided in Table 1(a). Here, the average constraint violation is computed over trajectories that hit the target with constraint violation. The constraint is computed by $({\text{center of obstacle}-\text{position of vehicle}})^{2}-({\text{radius ofobstacle}})^{2}.$ The results illustrate that SVDDP outperforms the other methods in all metrics. In addition, the results indicate that a proper choice of a margin of constraints increases the success rate of DDPs even further. The statistics of the computational time of one MPC call over ten different runs are presented in Table 2(a). The algorithms are implemented with JAX [33] and executed with Nvidia GeForce RTX 3080 GPU. Here, the prediction horizon amounts to 1.2 sec as in section E.1. Although SVDDP takes longer time than other methods, it is still short enough for real-time implementation. Finally, we emphasize that DDPs can generate smoother control trajectories as in Fig. E.3.
@@
Quadrotor: One of the obstacle fields and results are shown in Fig. 2(c) and Fig. 2(d). The obstacles are cylinders whose axis aligns with the $y$ axis with radii of 0.4. The initial and target points are [0, 0, 0] and [5, 0, 5]. The statistics of the results and computational time are provided in Table 1(b), and 2(b). The prediction horizon is 0.5 sec as in section E.2. Again, SVDDP outperforms other methods in success rate. It takes longer time for computation, but still fast enough for real-time application.
@@
Figure 2: Comparison of MPPIs and MPC DDPs for 2D Car and Quadrotor in one selected cluttered environment each. $\bm{\times}$:Initial positions. $\bigstar$:Target positions. : position at terminal time step.

### v2

# Second-Order Stein Variational Dynamic Optimization

@@ 1 Introduction
Trajectory Optimization (TO) is essential for robotic and autonomous systems to operate safely in dense and cluttered environments. From an optimization point of view, the planning and trajectory optimization problems in robotics have some unique characteristics. These include the non-convexity of the objective function in consideration, the nonlinear state and actuation constraints, and the highly nonlinear dynamics that characterize most robotic systems. The characteristic of non-convexity is due to the existence of multiple local minima. Such situations arise in reaching and navigation tasks performed in cluttered and dense environments.
There is a long history of TO methods and their application to robotics and autonomous systems. Starting from the Differential Dynamic Programming (DDP) [1] and the Iterative Pontryagin Maximum Principle (iPMP) [2], these optimal control methods have been extensively used in robotics. Both DDP and iPMP belong to the same family of algorithms that are gradient-based and rely on forward-backward sweeps. Their mathematical structure is derived using either the Bellman Principle of Optimality or the Pontryagin Maximum Principle and consists of the backpropagation of the value function or costate and forward propagation of state trajectories. This step of backpropagation is performed locally around the trajectories that are updated at each iteration. While the locality of the backward pass makes DDP and iPMP scalable and fast, it often results in sub-optimal solutions.
@@
In this paper, we investigate two mechanisms that combine sampling with gradient-based optimization and parallelism, and develop algorithms for planning and MPC with highly non-convex objectives. The first mechanism is a schedule of optimization steps, that consists of number of Netwon steps followed by an iteration that relies on sampling. The motivation for this hybrid mechanism is that while Newton steps can locally optimize the objective function, sampling can disrupt this process to ensure that optimization is not getting trap into a sub-optimal local minima. In the context of TO for robotics, this mechanism was first introduced in the work on Maximum Entropy DDP (MEDDP) and its unimodal and multimodal extensions. The difference of this work with respect to [17] is that here we show the benefits of this mechanisms in the context of MPC. It is also interesting to note that this hybrid optimization mechanism, which we name Messy Optimization, was also proposed in the area of business administration and management science [18] as mental model for enhances creativity and improving innovation.
The second mechanism goes one step further than MEDDP and results in new algorithm, the Stein Variational Differential Dynamic Programming (SVDDP). Grounded on the Stein Variational Newton’s Method (SVNM) [19], SVDDP is a kernel-based extension of MEDDP. In SVDDP, the optimization process alternates between two phases. Starting from different initial trajectories/particles, the first phase consists of multiple DDP updates performed in parallel fashion resulting in optimized state and control trajectories. In the second phase, these trajectories are updated only once with DDP using the SVNM to update the controls and generate the new state trajectories. This phase enforces exploration by pushing the trajectories to stay away from each other. In addition, the fact that this update is performed once every few iterations of the first phase ensures that the overall optimization will not be disrupted. We benchmark the proposed algorithm against variations on MEDDP and MPPI for both TO and MPC and show that it matches and, in most cases, outperforms these algorithms. The rest of the paper is organized as follows. Section 2 introduces the background of our work which includes MEDDP and SVGD and SVNM. In Section 3 we derive SVDDP. Section 4 includes experimental results, and section 5 concludes with future extensions.
@@ 2 Background
2.1 Maximum Entropy Dynamic Second-Order c Optimization
Consider a TO problem for a dynamical system with state $x\in\mathbb{R}^{n_{x}}$ and control $u\in\mathbb{R}^{n_{u}}$. Given the deterministic dynamics $x_{t+1}=f(x_{t},u_{t})$, The TO problem is formulated as follows:
@@ 2.2 Maximum Entropy DDP
where $Q(x,u)$ is the state-action value function ($Q$ function) of DDP and $Z(x)$ is the corresponding partition function. We refer readers to [17] and Appendix A for derivation. This indicates that the optimal policy has a form of unimodal Gaussian $\pi^{\ast}(\delta u|\delta x)\sim\mathcal{N}(\delta u^{\ast},\alpha Q_{uu}^{-1})$, where $\delta u^{\ast}$ is a solution of vanilla deterministic DDP:
$\delta{u}^{\ast}=\kappa+K\delta x,\quad\text{with}\ \kappa=-{Q}^{-1}_{{uu}}{Q_{{u}}},\ K=-{Q}^{-1}_{{uu}}{Q_{{ux}}}.$ (5)
@@ 2.3 Stein Variational Gradient Descend and Newton’s method
This section briefly covers Stein Variational Gradient Descent (SVGD) and SV Newton’s Method (SVNM). We first introduce the update law of SVGD, then describe the connection to functional gradient descent. Finally, we review SVNM as an extension to the second-order optimization of SVGD. For further information, we refer readers to [16] for SVGD and [22] for SVNM. In this section, we use $x\in\mathbb{R}^{d}$ to denote a variable of a function but not a state of a system. In addition, we use $x_{n}\in\mathbb{R}^{d}$ to denote a sampled point.
Stein Variational Gradient Descend: Let $p$ on $\mathbb{R}^{d}$ be a target distribution that we would like to approximate using a collection of samples. We drop the argument of the function $p$, and specify it only when we evaluate it at point $x_{n}$ as $p(x_{n})$. With samples $\{x_{n}\}$ from a tractable reference distribution $q$ on $\mathbb{R}^{d}$, SVGD iteratively computes a transport map $T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ so that the transformed samples from $q$, i.e., $\{T(x_{n})\}$ can be an empirical approximation of $p$. This map is obtained by solving the following optimization problem.
@@
where $l$ stands for $l$ th iteration. Here, $D_{\rm{KL}}$ is KL divergence that measures the difference between the two distributions. To simplify this problem, SVGD considers the vector-valued Reproducing Hilbert Kernel Space (RKHS) $\mathcal{H}^{d}=\mathcal{H}\times\cdots\times\mathcal{H}$, where $\mathcal{H}$ is a scalar-valued RKHS with kernel $k(x,x^{\prime})$. With RKHS, one can represent a function as a weighted composition of kernels, centered at samples. This representation also allows us to perform optimization in functional space by working with the weights in the same way as the normal optimization technique, such as GD. Here, the transportation map is written as a perturbation of the identity map denoted by $I$ along $\mathcal{H}^{d}$ as
@@
With this $T^{l}$, we define the objective of the problem (7) as
$\hat{J}_{q^{l}}[Q]=D_{\rm{KL}}\big{(}(I+Q)\ast q^{l}||p\big{)},$ (8)
which should be decreased by the proper choice of $Q$ from $\hat{J}_{q^{l}}[\bm{0}]$, where $\bm{0}(x)=0$ is a zero map. Since this $Q$ lies in the steepest direction of the functional $\hat{J}_{q}^{l}$ at $\bm{0}$, the map is given as follows.
@@
with a step size $\epsilon$. Observe the similarity with GD update of a normal optimization, i.e., $x^{l+1}=x^{l}+\epsilon\nabla f_{0}(x^{l})$ with objective $f_{0}$. The descent direction satisfies the following condition with $S\in\mathcal{H}^{d}$,
@@ 3.1 ME DDP with Functional Gradient and Hessian
We consider $N$ trajectories optimized by DDP that we would like to spread out for exploration. Using the deviation $u^{\ast}=\bar{u}+\delta u^{\ast}$, the gradient operator $\nabla_{u\ast}$ for $Q$ function of DDP is equivalent to $\nabla_{\delta u}$. Therefore, with the quadratically approximated $Q$, and the substitution of $\pi^{\ast}(u^{\ast})$ for $p(x)$ in the functional gradient (10), we get:
@@
where the derivatives of $Q$ are evaluated at $\bar{u}$. The first term on the right-hand side is zero because of the optimality condition used in the derivation of DDP. Therefore, we are left with the repulsive force terms, which seem attractive for spreading trajectories. However, since we lose the effect of $\alpha$, this formulation cannot capture the relative significance of the entropy in the objective. We need SVNM to capture it. By substituting $\pi^{\ast}(u)$ for $p(x)$ in the approximated Hessian (13) with the terms in (12), and performing empirical approximation, the Hessian for SV Newton’s method is obtained as:
$H^{s,s}=\frac{1}{N}\sum_{n=1}^{N}\Big{[}\frac{Q^{(n)}_{uu}}{\alpha}k(u^{(n)},u^{(s)})^{2}+\nabla_{u^{(n)}}k(u^{(n)},u^{(s)})\nabla_{u^{(n)}}k(u^{(n)},u^{(s)})^{\mathsf{T}}\Big{]},$ (15)
@@
$u^{(s)}\leftarrow u^{(s)}+\epsilon w(u^{(s)})+K\delta x,\quad\text{with}\quad w(u^{(s)})=\sum_{n=1}^{N}\beta^{(n)}k(u^{(n)},u^{(s)}),$ (16)
where $K$ is feedback gain of DDP in (A.10). Now, $\alpha$ is captured in $H^{s,s}$. Hence, we believe that this should be the correct formulation to incorporate the SV method with DDP. The second term in (15) is known as an expected outer product of the gradient, which captures information on the variation in all directions [23]. Thus, it can be seen as an approximated Hessian of the kernel. We also provide another example of interpretation in the Appendix B.2. To comprehend how the Hessian affects the repulsive force term, let us consider an example with $H\in\mathbb{R}^{2\times 2}$. We can draw an ellipsoid from its eigenvectors and eigenvalues whose major and minor axes point to the directions where $Q$ and $k$ increase most rapidly and slowly, respectively. When solving (13), and $H$ is inverted, the scaling factors of the axes, i.e., eigenvalues, are also inverted. However, the eigenvectors are preserved. Consequently, the axes are flipped. Hence, the major and minor axes of $H^{-1}$ indicate the direction in which $Q$ and $k$ increase most slowly and rapidly, respectively. When (13) is solved, the repulsive force term is projected in the direction where $Q$ and $k$ increases slowly. This means that the force is transformed so that it does not disturb the optimization process too much and spreads trajectories more by the curvature information on $Q$ and $k$, respectively. This projection leads to an efficient exploration. We provide the schematic of SVDDP in Fig.1. We note that the update (16) is not applied to the best trajectory, ensuring that at least one mode shows a monotonic cost reduction to preserve the convergence property of DDP as in [17, 24]. We note that we use cost as a metric, although other metrics e.g., diversity, can be used depending on the purpose.
@@ 4 Experimental Results
This section provides experimental results, including TO and MPC. For TO, we compare normal DDP, Unimodal, Multimodal Gaussian MEDDPs (UG-, MG-MEDDP), and SVDDP. We note that DDPs are modified so that they can handle constraints as in section 3.3. For MPC, our comparisons involve DDP-MPC, UG-MEDDP, MG-MEDDP, SVDDP in MPC mode. In addition, MPPIs with three different policy parameterization and update laws are included in the comparison. They are Unimodal Gaussian MPPI (UG-MPPI), Multimodal Gaussian MPPI (MG-MPPI), and Unimodal Gaussian with Stein Variational update law (SV-MPPI) [7]. These three MPPIs correspond to the UG-MEDDP, MG-MEDDP, and SVDDP, respectively. In terms of systems, we use three systems, including a 2D Car, quadrotor, and 7DoF arm in TO and 2D car and quadrotor for MPC. In addition, we have comparisons of DDPs in swimmer and ant dynamics in Brax [31] whose results are presented in section F.
Trajectory Optimization: This experiment is to examine the exploration capability and convergence of SV and ME DDPs over normal DDP. In addition, we observe the effect of temperature $\alpha$, number of modes, and stepsize $\epsilon$ on the performance. Figure D.1 provides an overview of the experiments. We performed the same experiment 15 times to capture the properties of the algorithms, and 15 outcomes are shown in the figure. The task is to reach the target while avoiding obstacles under control limits. The configuration and shapes of the obstacles give multiple local minima, which normal DDP without exploration gets captured by poor ones. ME and SVDDPs can explore and find better solutions with the exploration mechanisms. To examine the effect of temperature $\alpha$ and the number of modes (trajectories), we performed the same experiments, with different values of them. The results of the experiments are provided in Table D.1. For SVDDP, we also changed step size $\epsilon$s. We plot the evolution of the mean cost over iteration in Fig. D.2 to compare the convergence speed of DDPs.
@@
Model Predictive Control: We perform experiments in MPC with the 2D car and quadrotor system used in TO for similar reaching tasks. With 2D car and quadrotor, we compare seven algorithms of MPC DDPs and one of the most successful sampling-based algorithms MPPI to examine the performance of our algorithms. We have DDP, UG-MEDDP, MG-MEDDP, and SVDDP in MPC mode for DDP-based algorithms. For MPPI, we have UG-, MG-, and SV-MPPI. Details of the experiments are provided in Appendix E. We use [32] to tune MPPIs parameters such as covariance and temperature, etc.
@@
In the experiments, a trajectory is considered successful when it reaches a ball centered at the target with radius 0.5, and does not hit the obstacles. The experiments are performed ten times per algorithm and environment, except for the deterministic normal DDP. This means that for each algorithm we will have 100 runs from which we collect statistics on the performance. To show that DDP is the effective method for ME second-order optimization, we also perform the same experiment by discarding the feedback gain $K$ in (5) and (16) when sampling trajectories in ME and SVDDPs. This is because providing feedback gain is a unique property of DDP in contrast to other second-order methods.
@@
2D Car: One of the obstacle fields and results are shown in Fig. 2(a) and Fig. 2(b). The initial point and the target are [0, 0] and [5, 5]. The normal DDP get stack while MEDDPs and SVDDPs are able to hit the target. MPPIs can explore the fields, but not as good as ME and SV DDPs and they tend to require longer time even when they escape local minima. Statistics on the performance of the algorithms with respect to the success rate are provided in Table 1(a). Here, the average constraint violation is computed over trajectories that hit the target with constraint violation. The constraint is computed by $({\text{center of obstacle}-\text{position of vehicle}})^{2}-({\text{radius ofobstacle}})^{2}.$ The results illustrate that SVDDP outperforms the other methods in all metrics. In addition, the results indicate that a proper choice of a margin of constraints increases the success rate of DDPs even further. Without feedback gains in sampling, the algorithms reduce the success rate as SVDDP: 74 %, MG-MEDDP: 71 % and UG-MEDDP 68 %, which highlights the importance of feedback gains when sampling trajectories. The statistics of the computational time of one MPC call over ten different runs are presented in Table 2(a). The algorithms are implemented with JAX [33] and executed with Nvidia GeForce RTX 3080 GPU. Here, the prediction horizon amounts to 1.2 sec as in section E.1. Although SVDDP takes longer time than other methods, it is still short enough for real-time implementation. Finally, we emphasize that DDPs can generate smoother control trajectories as presented in Fig. E.3, which shows one of the two control commands with control bounds.
@@
Quadrotor: One of the obstacle fields and results are shown in Fig. 2(c) and Fig. 2(d). The obstacles are cylinders whose axes aligns with the $y$ axis. The initial and target points are [0, 0, 0] and [5, 0, 5]. The statistics of the results and computational time are provided in Table 1(b), and 2(b). Again, SVDDP outperforms other methods in success rate. It takes longer for computation, but still fast enough for real-time application. Without feedback gains in sampling, we obtained the following success rate. SVDDP: 45 %, MG-MEDDP: 56 % and UG-MEDDP 52 %, which indicates the importance of the feedback gain even more in an unstable system.
@@
Figure 2: Comparison of MPPIs and MPC DDPs for 2D Car and Quadrotor in one of the ten obstacle fields. $\bm{\times}$:Initial positions. $\bigstar$:Target positions. : position at terminal time step.